OpenStudy (anonymous):

Solving systems of linear equations using substitution.... y=x-3; x+y=5

OpenStudy (anonymous):

use substitution to isolate one variable (already done here) and use it to solve the other x+(x-3) = 5 x + x = 8 2x = 8 x = 4 use that to solve first half y = 4 - 3 y = 1

OpenStudy (anonymous):

2x - 3 = 5 x = 4 sub in isr equn, y = 4-3 = 1 x=4, y=1

OpenStudy (anonymous):

How did you come up with 8?

OpenStudy (anonymous):

add three to both sides

OpenStudy (anonymous):

you need to solve for x so you need x by itself

OpenStudy (anonymous):

you could do this the other way around and solve for y first too, I just used x first

OpenStudy (anonymous):

Thank you, that is very helpful

OpenStudy (anonymous):

yw